# Continuous-time Risk-sensitive Reinforcement Learning via Quadratic Variation Penalty

### v1

@@ Abstract
This paper studies continuous-time risk-sensitive reinforcement learning (RL) under the entropy-regularized, exploratory diffusion process formulation with the exponential-form objective. The risk-sensitive objective arises either as the agent’s risk attitude or as a distributionally robust approach against the model uncertainty. Owing to the martingale perspective in Jia and Zhou, (2023), the risk-sensitive RL problem is shown to be equivalent to ensuring the martingale property of a process involving both the value function and the q-function, augmented by an additional penalty term: the quadratic variation of the value process, capturing the variability of the value-to-go along the trajectory. This characterization allows for the straightforward adaptation of existing RL algorithms developed for non-risk-sensitive scenarios to incorporate risk sensitivity by adding the realized variance of the value process. Additionally, I highlight that the conventional policy gradient representation is inadequate for risk-sensitive problems due to the nonlinear nature of quadratic variation; however, q-learning offers a solution and extends to infinite horizon settings. Finally, I prove the convergence of the proposed algorithm for Merton’s investment problem and quantify the impact of temperature parameter on the behavior of the learning procedure. I also conduct simulation experiments to demonstrate how risk-sensitive RL improves the finite-sample performance in the linear-quadratic control problem.
@@ 1 Introduction
Continuous-time reinforcement learning (RL) has drawn considerable attention owing to its practical significance in modeling systems that necessitate or benefit from high-frequency or real-time interaction with the environment, such as in financial trading (Wang et al., 2023a, ), real-time response systems (Andersson et al., , 2015), and robotics (Kober et al., , 2013). Furthermore, its theoretical development serves to bridge the gap between the conventional stochastic control and the discrete-time, Markov decision process (MDP)-based RL theory.
Existing literature on RL largely focuses on optimizing the expectation of a flow of time-additive rewards. Here, the instantaneous reward can be interpreted as a utility within the time-additive von Neumann–Morgenstern expected utility framework. Though utility functions can reflect the agent’s attitude towards risk in a one-off setting, economic literature has long recognized that such an additive form inadequately captures intertemporal preferences on the uncertainty (Epstein and Zin, , 1989). Moreover, empirically measuring and incorporating an agent’s risk preference via a simple additive utility function is notorious difficult.
An alternative approach is the risk-sensitive control (which can be traced back to Jacobson, 1973), and it later becomes popular, particularly in financial asset management, e.g., in Bielecki and Pliska, (1999). In contrast to solely considering the expectation, risk-sensitive objective accounts for the whole distribution of the accumulated reward. Moreover, the risk-sensitive objective function in the exponential form is well-known to be closely related to the robustness within a family of distributions measured by the Kullback–Leibler (KL) divergence, which is also known as the robust control problems (Hansen and Sargent, , 2001). Such uncertainty on the distribution of a random variable (not just its realization whose uncertainty can be statistically quantified) is often regarded as the Knightian uncertainty or ambiguity, which often occurs due to the lack of knowledge or historical data (LeRoy and Singell Jr, , 1987). From this perspective, an RL agent often encounters the similar situation, in which the agent lacks information about the environment, and hence, has difficulty formulating a probabilistic model to quantify the associated risk. Therefore, it is natural to consider the risk sensitivity in the RL setting.
@@
The primary contribution of this paper lies in the establishment of the q-learning theory for the continuous-time risk-sensitive RL problems. The conventional Q-function for the discrete-time, MDP-based risk-sensitive RL often relies nonlinear recursive relations, such as “exponential Bellman equation” (Fei et al., , 2021) or the “distributional robust Bellman equation” (Wang et al., 2023b, ). However, in continuous-time scenarios, the complicated structure gets simplified and clarified. This simplification stems from two key observations: firstly, a risk-sensitive control problem can be equivalently transformed into its non-risk-sensitive counterpart augmented with a quadratic variation (QV) penalty (Skiadas, , 2003); secondly, the conventional Q-function has to be properly decomposed and rescaled into the value function and q-function in continuous time (Jia and Zhou, , 2023), and the latter does not have nonlinear effect. In particular, the definition and characterization of the risk-sensitive q-function is almost parallel to that for the non-risk-sensitive q-function established in Jia and Zhou, (2023), and they only differ by an extra term involving the QV of the value-to-go process (the value function applying on the state variables). Consequently, explicit computation of exponential forms becomes unnecessary, and the derived martingale condition is linear in the risk-sensitive q-function, albeit still nonlinear in the value function. This linearity facilitates the application of algorithms aimed at enforcing the martingale property of a process, as discussed in Jia and Zhou, 2022a , to risk-sensitive RL problems.
@@
The second contribution is to analyze the proposed RL algorithm in Merton’s investment problem with power utility, which can be viewed as the risk-sensitive objective of the log-return of the portfolio. This problem has been solved only in theory, in which the model for stock price is given and known. In contrast, I study how agent can learn the policy with no prior knowledge. Specifically, I investigate the role of the temperature parameter in the entropy-regularized RL. While many have suggested intuitively that the temperature parameter governs the tradeoff between exploitation and exploration, its algorithmic impact has yet to be formally studied. To the author’s best knowledge, there is no clear guidelines on how to choose the temperature parameter endogenously, except for some theoretical results on the convergence or asymptotic expansion of the exploratory problem to the classical problem in the non-risk-sensitive setting, e.g., Wang et al., (2020); Tang et al., (2022); Dai et al., 2023b . These papers all consider the difference between the difference between the total reward under the optimal deterministic/stochastic policy by ignoring the estimation error of optimal policy incurred in a specific learning procedure. Therefore, from their analysis, it seems that randomization and entropy regularization always cause inefficiency but have no clear benefits.
I highlight that the virtue of entropy regularization lies on the algorithmic aspect via boosting the estimation accuracy of the optimal policy. Within the framework of stochastic approximation algorithms, this study reveals that the temperature parameter functions analogously to the learning rate: Higher temperatures correspond to faster learning because data collected spans larger space and contains more information (interpreted as exploration) but also entail higher noise in the data and lower expected reward (due to a lack of exploitation). The convergence and convergence rate are determined by the combined schedule of the step size and the temperature parameter in each iteration. I give two possible conditions to ensure the proposed learning algorithm converges at the optimal rate that matches the state-of-the-art in terms of the number of episodes.
@@
Previous literature on risk-sensitive control problems in continuous time studies their theoretical properties and economic implications, for example, the well-posedness (Fleming and McEneaney, , 1995; Nagai, , 1996; Dupuis et al., , 2000; Fleming and Sheu, , 2002; Menaldi and Robin, , 2005), and how risk sensitivity explains the equity premium puzzle (Maenhout, , 2004; Glasserman and Xu, , 2013). However, in the conventional paradigm, agents have risk-sensitive objectives, but they know a benchmark distribution (or a model) and also know the forms of the reward functions. RL perspective contrasts the conventional settings in not knowing a benchmark distribution (i.e., being model-free), and even not the forms of reward functions. The latter may not be a concern for applications in trading and pricing as the reward is simply the return or profit, whereas could be critical in general RL context in which the agent has to execute a certain action to figure out the associated reward. For the former, one way to specify the benchmark distribution is to fit a simple model within a parametric family, e.g., geometric Brownian motions. However, due to the restriction of the KL divergence, this approach implicitly assumes that the ground truth model is absolutely continuous with respect to the estimated model. This could still wrongly specify a family of models. Moreover, RL is sometimes conducted in an online fashion, where no past data is available and data needs to be collected as part of the solution. In such scenarios, it is even more difficult to specify a benchmark model.
@@
The continuous-time RL framework adopted in this paper is developed [^1] It is worthwhile pointing out that prior literature on continuous-time RL has been largely restricted to deterministic system, such as Baird, (1994); Doya, (2000); Tallec et al., (2019); Lee and Sutton, (2021); Kim et al., (2021). To the author’s knowledge, Wang et al., (2020) is the first paper to approach stochastic control problems from an RL perspective rather than a computational method. in Wang et al., (2020), who introduce stochastic policies and entropy regularization to the standard stochastic control problems to formulate RL in continuous-time, and to embed RL into relaxed control problems (Fleming and Nisio, , 1984; Zhou, , 1992) for theoretical analysis. Moving from the theoretical framework to general principles to devise data-driven RL algorithms, Jia and Zhou, 2022a propose the unifying martingale perspective in the policy evaluation stage and suggest the loss function and orthogonality conditions to approximate the martingale properties. Jia and Zhou, 2022b further derive the representation for the policy gradient for stochastic policies. Moreover, Jia and Zhou, (2023) clarify the notion of “q-function” in the continuous-time system, and reveal the relation between the q-function in RL and Hamiltonian in control literature. In particular, Jia and Zhou, (2023) show that the q-function and the value function can be learned jointly via martingale conditions, thus, the principles proposed in Jia and Zhou, 2022a can be applied to devise q-learning algorithms. Based on this framework, there have been extensive studies and applications, such as Guo et al., (2022); Frikha et al., (2023); Wei and Yu, (2023) for mean-field interactions, Dai et al., 2023a for time-inconsistency, Wang and Zhou, (2020); Szpruch et al., (2024) for linear-quadratic controls, and Wang et al., 2023a for optimal execution. However, all the above has been done only for the ordinary, non-risk-sensitive RL. This paper focuses on the risk-sensitive objective, to see how the martingale perspective can be adopted and applied here and how to design algorithms based on it.
The important observation on the link between a robust control problem and the QV penalty is made by Skiadas, (2003). Note that the risk-sensitive objective in the exponential form naturally connects to the KL divergence that is used in the robust control problems, and the exponential martingale is closely related to the QV process. Thus, a risk-sensitive RL problem can also be transformed to an ordinary, non-risk-sensitive RL problem plus an extra QV penalty on the value function along the trajectory. Therefore, it becomes a stochastic differential utility or recursive utility maximization problem (Duffie and Epstein, , 1992), and hence, the martingale optimality principle still applies to characterize the optimal policy. Based on the framework in Jia and Zhou, (2023), I develop a martingale characterize of the optimal policy and the optimal value function, also involving the QV penalty term. Consequently, q-learning algorithms can be designed to approximate martingale processes based on ideas in Jia and Zhou, 2022a . Recall that, despite that q-function is defined as the combination of the value function’s up to second-order derivatives and nonlinear terms, q-function can be approximated as a whole independently of the value function. Thus, the high-order derivatives and nonlinear terms do no cause difficulty.
@@
The same risk-sensitive objective in the exponential form has also been studied in the discrete-time, tabular MDP-based RL settings, e.g., Borkar, (2002); Fei et al., (2020, 2021); Wang et al., 2023b . The approaches adopted are via “exponential Bellman equation” or “distributional robust Bellman equation”, both of which involve nonlinear recursive relation. The best regret bound in terms of the number of episodes is at the square-root order, ignoring the logarithm factor. But these works do not include entropy regularizations. In the continuous-time, the resulting martingale condition is still nonlinear in the value function while becomes linear in q-function. Numerically, it avoids computation of exponential forms and reduces the sampling noises. In the Merton’s investment problem, the same convergence rate in terms of the number of episodes, matching the state-of-the-art, can also be proved when the underlying stock price follows the geometric Brownian motion.
@@ 2.1 Classical model-based formulation
Second, the form of the objective function (2) can be viewed as applying an exponential utility function on the total reward, where $-\epsilon$ measures the absolute risk aversion. Whereas, the risk sensitivity should not be confused with the commonly used notion of risk-aversion in the additive utility functions in economics literature, in which the risk-aversion may be represented via the concavity of the function $r$ and $h$.
Third, the risk sensitivity also stems from the ambiguity and robustness in the sense of Hansen and Sargent, (2011), or sometimes called the control under model misspecification (Hansen and Sargent, , 2001) or robust control (Dupuis et al., , 2000). From this perspective, the agent is still maximizing the expected total reward whereas is uncertain about the true distribution, hence the agent forms “multi-prior” in the sense of Gilboa and Schmeidler, (1989) and considers the worse distribution:
@@
    $\int_{t}^{s}\left\{r(u,X_{u}^{\boldsymbol{a}^{*}},a_{u}^{*})\mathrm{d}u+\frac{\epsilon}{2}\mathrm{d}\langle V^{*^{\boldsymbol{a}^{*}}}\rangle(u)\right\}+V^{*}(s,X_{s}^{\boldsymbol{a}^{*}};\epsilon)$ is an ($(\{\mathcal{F}_{s}^{W}\}_{s\geq t},\mathbb{P}^{W})$- (local) martingale; and
@@
    for any $\boldsymbol{a}$, $\int_{t}^{s}\left\{r(u,X_{u}^{\boldsymbol{a}},a_{u})\mathrm{d}u+\frac{\epsilon}{2}\mathrm{d}\langle V^{*^{\boldsymbol{a}}}\rangle(u)\right\}+V^{*}(s,X_{s}^{\boldsymbol{a}};\epsilon)$ is an ($(\{\mathcal{F}_{s}^{W}\}_{s\geq t},\mathbb{P}^{W})$- (local) supermartingale,

### v2

@@ Abstract
This paper studies continuous-time risk-sensitive reinforcement learning (RL) under the entropy-regularized, exploratory diffusion process formulation with the exponential-form objective. The risk-sensitive objective arises either as the agent’s risk attitude or as a distributionally robust approach against the model uncertainty. Owing to the martingale perspective in Jia and Zhou (J Mach Learn Res 24(161):1–61, 2023), the risk-sensitive RL problem is shown to be equivalent to ensuring the martingale property of a process involving both the value function and the q-function, augmented by an additional penalty term: the quadratic variation of the value process, capturing the variability of the value-to-go along the trajectory. This characterization allows for the straightforward adaptation of existing RL algorithms developed for non-risk-sensitive scenarios to incorporate risk sensitivity by adding the realized variance of the value process. Additionally, I highlight that the conventional policy gradient representation is inadequate for risk-sensitive problems due to the nonlinear nature of quadratic variation; however, q-learning offers a solution and extends to infinite horizon settings. Finally, I prove the convergence of the proposed algorithm for Merton’s investment problem and quantify the impact of temperature parameter on the behavior of the learning procedure. I also conduct simulation experiments to demonstrate how risk-sensitive RL improves the finite-sample performance in the linear-quadratic control problem.
@@ 1 Introduction
Continuous-time reinforcement learning (RL) has drawn considerable attention owing to its practical significance in modeling systems that necessitate or benefit from high-frequency or real-time interaction with the environment, such as in financial trading, real-time response systems, and robotics. Furthermore, its theoretical development serves to bridge the gap between the conventional stochastic control and the discrete-time, Markov decision process (MDP)-based RL theory.
Existing literature on RL largely focuses on optimizing the expectation of a flow of time-additive rewards. Here, the instantaneous reward can be interpreted as a utility within the time-additive von Neumann–Morgenstern expected utility framework. Though utility functions can reflect the agent’s attitude towards risk in a one-off setting, economic literature has long recognized that such an additive form inadequately captures intertemporal preferences on the uncertainty (Epstein and Zin, , 1989). Moreover, empirically measuring and incorporating an agent’s risk preference via a simple additive utility function is notoriously difficult.
An alternative approach is the risk-sensitive control (which can be traced back to (Jacobson, , 1973)), and it later becomes popular, particularly in financial asset management, e.g., in Bielecki and Pliska, (1999). In contrast to solely considering the expectation, risk-sensitive objective accounts for the whole distribution of the accumulated reward. Moreover, the risk-sensitive objective function in the exponential form is well-known to be closely related to the robustness within a family of distributions measured by the Kullback–Leibler (KL) divergence, which is also known as the robust control problems (Hansen and Sargent, , 2001). Such uncertainty on the distribution of a random variable (not just its realization whose uncertainty can be statistically quantified) is often regarded as the Knightian uncertainty or ambiguity, which often occurs due to the lack of knowledge or historical data (LeRoy and Singell Jr, , 1987). From this perspective, an RL agent often encounters the similar situation, in which the agent lacks information about the environment, and hence, has difficulty formulating a probabilistic model to quantify the associated risk. Therefore, it is natural to consider the risk sensitivity in the RL setting.
@@
The primary contribution of this paper lies in the establishment of the q-learning theory for the continuous-time risk-sensitive RL problems. The conventional Q-function for the discrete-time, MDP-based risk-sensitive RL often relies on nonlinear recursive relations, such as “exponential Bellman equation” (Fei et al., , 2021) or the “distributional robust Bellman equation” (Wang et al., 2023b, ). However, in continuous-time scenarios, the complicated structure gets simplified and clarified. This simplification stems from two key observations: firstly, a risk-sensitive control problem can be equivalently transformed into its non-risk-sensitive counterpart augmented with a quadratic variation (QV) penalty (Skiadas, , 2003); secondly, the conventional Q-function has to be properly decomposed and rescaled into the value function and q-function in continuous time (Jia and Zhou, , 2023), and the latter does not have nonlinear effect. In particular, the definition and characterization of the risk-sensitive q-function is almost parallel to that for the non-risk-sensitive q-function established in Jia and Zhou, (2023), and they only differ by an extra term involving the QV of the value-to-go process (the value function applying on the state variables). Consequently, explicit computation of exponential forms becomes unnecessary, and the derived martingale condition is linear in the risk-sensitive q-function, albeit still nonlinear in the value function. This linearity facilitates the application of algorithms aimed at enforcing the martingale property of a process, as discussed in Jia and Zhou, 2022a , to risk-sensitive RL problems.
@@
The second contribution is to analyze the proposed RL algorithm in Merton’s investment problem with power utility, which can be viewed as the risk-sensitive objective of the log-return of the portfolio. This problem has been solved only in theory, in which the model for stock price is given and known. By contrast, I study how an agent can learn the policy with no prior knowledge. Specifically, I investigate the role of the temperature parameter in the entropy-regularized RL. While many have suggested intuitively that the temperature parameter governs the tradeoff between exploitation and exploration, its algorithmic impact has yet to be formally studied. To the author’s best knowledge, there are no clear guidelines on how to choose the temperature parameter endogenously, except for some theoretical results on the convergence or asymptotic expansion of the exploratory problem to the classical problem in the non-risk-sensitive setting, e.g., Wang et al., (2020); Tang et al., (2022); Dai et al., (2026). These papers all consider the difference between the total reward under the optimal deterministic/stochastic policy by ignoring the estimation error of optimal policy incurred in a specific learning procedure. Therefore, from their analysis, it seems that randomization and entropy regularization always cause inefficiency but have no clear benefits.
I highlight that the virtue of entropy regularization lies in the algorithmic aspect via boosting the estimation accuracy of the optimal policy. Within the framework of stochastic approximation algorithms, this study reveals that the temperature parameter functions analogously to the learning rate: Higher temperatures correspond to faster learning because data collected spans larger space and contains more information (interpreted as exploration) but also entail higher noise in the data and lower expected reward (due to a lack of exploitation). The convergence and convergence rate are determined by the combined schedule of the step size and the temperature parameter in each iteration. I give two possible conditions to ensure the proposed learning algorithm converges at the optimal rate that matches the state-of-the-art in terms of the number of episodes.
@@
Previous literature on risk-sensitive control problems in continuous time studies their theoretical properties and economic implications, for example, the well-posedness (Fleming and McEneaney, , 1995; Nagai, , 1996; Dupuis et al., , 2000; Fleming and Sheu, , 2002; Menaldi and Robin, , 2005), and how risk sensitivity explains the equity premium puzzle (Maenhout, , 2004; Glasserman and Xu, , 2013). However, in the conventional paradigm, agents have risk-sensitive objectives, but they know a benchmark distribution (or a model) and also know the forms of the reward functions. RL perspective contrasts the conventional settings in not knowing a benchmark distribution (i.e., being model-free), and even not the forms of reward functions. The important observation on the link between a risk-sensitive problem and the QV penalty is made by Skiadas, (2003). Thus, a risk-sensitive RL problem can also be transformed to an ordinary, non-risk-sensitive RL problem plus an extra QV penalty on the value function along the trajectory. Therefore, it becomes a recursive utility maximization problem (Duffie and Epstein, , 1992), and hence, the martingale optimality principle still applies to characterize the optimal policy.
@@
The continuous-time RL framework adopted in this paper is developed [^1] It is worthwhile pointing out that prior literature on continuous-time RL has been largely restricted to deterministic system, such as Doya, (2000); Lee and Sutton, (2021); Kim et al., (2021) and many others. To the author’s knowledge, Wang et al., (2020) is the first paper to approach stochastic control problems from an RL perspective rather than a computational method. in Wang et al., (2020), who introduce stochastic policies and entropy regularization to the standard stochastic control problems to formulate RL in continuous-time, and to embed RL into relaxed control problems for theoretical analysis. Moving from the theoretical framework to general principles to devise data-driven RL algorithms, Jia and Zhou, 2022a propose the unifying martingale perspective in the policy evaluation stage and suggest the loss function and orthogonality conditions to approximate the martingale properties. Jia and Zhou, 2022b further derive the representation for the policy gradient for stochastic policies. Moreover, Jia and Zhou, (2023) clarify the notion of “q-function” in the continuous-time system, and reveal the relation between the q-function in RL and Hamiltonian in control literature. In particular, Jia and Zhou, (2023) show that the q-function and the value function can be learned jointly via martingale conditions, thus, the principles proposed in Jia and Zhou, 2022a can be applied to devise q-learning algorithms. Based on this framework, there have been extensive studies and applications, such as Guo et al., (2022); Frikha et al., (2025); Wei and Yu, (2025) for mean-field interactions, Dai et al., (2023) for time-inconsistency, Wang and Zhou, (2020); Szpruch et al., (2024) for linear-quadratic controls, and Wang et al., 2023a for optimal execution. However, all the above has been done only for the ordinary, non-risk-sensitive RL. This paper focuses on the risk-sensitive objective, to see how the martingale perspective can be adopted and applied here and how to design algorithms based on it.
@@
The discrete-time counterparts (by replacing integrals with sums, and replacing controlled diffusion processes by MDPs, but without entropy regularization) are also popular in discrete-time robust RL, such as Borkar, (2002); Fei et al., (2020, 2021); Wang et al., 2023b ; Xu et al., (2025). On the other hand, the entropy-regularized RL, such as soft-Q-learning (Schulman et al., , 2017) and soft-actor-critic (Haarnoja et al., , 2018), and also has been integrated to continuous-time RL by Wang et al., (2020); Guo et al., (2022); Jia and Zhou, 2022b ; Jia and Zhou, (2023), and many others. The combination of risk-sensitive RL and entropy regularization has been adopted in discrete-time setting in Enders et al., (2024), but is novel to continuous-time setting. In the discrete-time, the analysis has been largely restricted to tabular cases (Borkar, , 2002; Fei et al., , 2020, 2021; Wang et al., 2023b, ), and their approaches are via “exponential Bellman equation” or “distributional robust Bellman equation”, both of which involve nonlinear recursive relation. The best regret bound in terms of the number of episodes is at the square-root order in general, ignoring the logarithm factor. But these works do not include entropy regularizations. By considering the problem in continuous-time, the resulting martingale condition is still nonlinear in the value function while becomes linear in q-function. Numerically, it avoids computation of exponential forms and reduces the sampling noises. Intuitively speaking, this simplification is due to the validity of the normal approximation for the state transition. In the Merton’s investment problem, the same convergence rate in terms of the number of episodes, matching the state-of-the-art, can also be proved when the underlying stock price follows the geometric Brownian motion.
The other example considered in this paper is the off-policy linear-quadratic control. The linear-quadratic control problems have been widely studied in the RL context, in both discrete-time (e.g., Hambly et al., (2021); Fazel et al., (2018); Wang et al., (2021); Abbasi-Yadkori and Szepesvári, (2011)) and continuous-time (e.g., Basei et al., (2022); Guo et al., (2023); Giegrich et al., (2024); Szpruch et al., (2024)). All these literature studies non-risk-sensitive problems. The only exception is Xu et al., (2025), who also study risk-sensitive problem but focus on online, on-policy learning by developing a model-based, least-square-type algorithm. The example considered in this paper has three distinctions compared to previous literature: first, the volatility of the noise in the system is controlled; second, it is the first paper considering the continuous-time risk-sensitive RL, and an entropy regularizer is included in the objective function; third, the learning is conducted off-policy. The formulation and the risk-sensitive q-learning theory introduced in this paper gives full flexibility and account for all these unconventional features compared to what has been studied in the literature, and provides the foundation for devising model-free learning algorithms.
@@ 2.1 Classical model-based formulation
Second, the form of the objective function (2) can be viewed as applying an exponential utility function on the total reward, where $\epsilon$ measures the absolute risk aversion. Whereas, the risk sensitivity should not be confused with the commonly used notion of risk-aversion in the additive utility functions in economics literature, in which the risk-aversion may be represented via the concavity of the function $r$ and $h$.
Third, the risk sensitivity also stems from the ambiguity and robustness in the sense of Hansen and Sargent, (2011), or sometimes called the control under model misspecification (Hansen and Sargent, , 2001) or robust control (Dupuis et al., , 2000). From this perspective, the agent is still maximizing the expected total reward whereas the agent is uncertain about the true distribution, hence the agent forms “multi-prior” in the sense of Gilboa and Schmeidler, (1989) and considers the worse distribution:
@@
    $\int_{t}^{s}\left\{r(u,X_{u}^{\boldsymbol{a}^{*}},a_{u}^{*})\mathrm{d}u+\frac{\epsilon}{2}\mathrm{d}\langle V^{*^{\boldsymbol{a}^{*}}}\rangle(u)\right\}+V^{*}(s,X_{s}^{\boldsymbol{a}^{*}};\epsilon)$ is an $(\{\mathcal{F}_{s}^{W}\}_{s\geq t},\mathbb{P}^{W})$- (local) martingale; and
@@
    for any $\boldsymbol{a}$, $\int_{t}^{s}\left\{r(u,X_{u}^{\boldsymbol{a}},a_{u})\mathrm{d}u+\frac{\epsilon}{2}\mathrm{d}\langle V^{*^{\boldsymbol{a}}}\rangle(u)\right\}+V^{*}(s,X_{s}^{\boldsymbol{a}};\epsilon)$ is an $(\{\mathcal{F}_{s}^{W}\}_{s\geq t},\mathbb{P}^{W})$- (local) supermartingale,